# Spectral Criteria for the Asymptotics of Local Functionals of Gaussian Fields and Their Application to Nodal Volumes

### v2

@@ Abstract
In this paper, we establish a general criterion for the positivity of the variance of the second chaotic component of local functionals of stationary vector-valued Gaussian fields. This criterion is formulated in terms of the spectral properties of the covariance function, without requiring assumptions of integrability or isotropy. It offers a simple and robust framework for analyzing variance asymptotics in such models. We apply this approach to the study of the nodal volume and the number of critical points of a Gaussian field, proving the positivity of the limiting variance under mild conditions on the covariance function. Additionally, we examine the asymptotics of the particular model of Euclidean random waves, deriving the central limit theorem through an analysis of the second and fourth chaotic components, thereby resolving an open question in large dimensions. As a byproduct, we condense and generalize many existing results on the volume of intersections of random waves, bypassing the need for traditional, intricate variance computations. Our findings shed new light on the second-chaos cancellation phenomenon from a spectral perspective and can be extended to any local, possibly singular, functional of Gaussian fields.
@@ 1.1 A general framework
Let $V$ and $W$ be two finite dimensional vector spaces of dimension $d$ and $n$ respectively. Let $X:V\rightarrow W$ be a vector-valued stationary Gaussian random field with a continuous non-degenerate covariance function $\Omega:V\rightarrow S^{2}(W^{*})$, where this last space denotes the set of symmetric bilinear map on $W$. We denote by $\eta$ the associated Gaussian measure on $W$, and let $f:W\rightarrow\mathbb{R}$ in $L^{2}(\eta)$. Our primal object of interest is the quantity
@@
A proof of the CLT is usually split into two very distinct problems: the asymptotic normality and the positivity of the limiting variance. The asymptotic normality is well-understood and was first proved in [Cuz76] in the case $d=n=1$ and $\Omega\in L^{1}(\mathbb{R}^{d})$ by approximation using $m$-dependent processes. Since then, it has been refined with the introduction of the Wiener chaos expansion, which has become an indispensable tool for proving asymptotics related to $Z_{\lambda}(\phi)$. In particular, the celebrated Breuer–Major theorem [Bre83] states that asymptotic normality holds if the Hermite rank of $f$ is $q$ and $\Omega\in L^{q}(V)$. Modern proofs of this theorem now include quantitative CLT and functional convergence [Nou11, Nou20, Nua20] and make use of more advanced techniques, such as Malliavin calculus and Stein’s method. The literature on the topic is extensive.
@@
thus the limiting variance is positive if and only if $\rho(0)$ is positive. In a more general framework, the variance can be understood from local properties of the spectral measure $\rho$ near the origin. To explicitly compute the measure $\rho$ in terms of the spectral measure $\mu$ associated to $X$, the most direct path is to expand $f$ into its orthogonal Hermite decomposition on the space $L^{2}(\eta)$. The variance will then be expressed as sum of powers of the covariance function $\Omega$. Passing to Fourier transform, this yields an expression of $\rho$ in terms of a sum of non-negative terms, given by iterated convolutions of the spectral measure $\mu$ and Hermite expansion of $f$. Additional comments on the advantage of this approach will be made after Theorem 1.4.
@@ 1.1.1 A general spectral expression for the variance
This corollary, which is a direct corollary of Theorem 1.1 and Lemma 3.8, puts forward the interaction between the iterated convolutions of the spectral measure, their regularity near the origin, and asymptotic results of local functionals of the Gaussian field. In the case $n=1$, the $q$-fold convolution of $\mu$ can be seen as the distribution of a random walk with transition probability distribution $\mu$ after $q$ steps. The properties of $\mu^{*q}$ near the origin are linked to the probability of return around 00 of the random walk. This analogy has been successfully used in [Lac22] to obtain variance asymptotics related to the volume of random excursions sets, and in [Gro24] to study the covariance structure of the Berry random wave model.
@@ 1.1.2 A general cancellation theorem for the variance
    If $q$ is even and $\operatorname{Im}(\mu^{*q/2})\subset C(f_{q})$, there is a competition between the cancellation phenomenon and the higher growth of the variance explained in the first two points. From Theorem 1.4, if the inclusion is sufficiently "smooth" then the $q$-th chaos cancels. It might not be the case for more irregular measures, where one can only rely on the general Corollary 1.3. The random waves model, whose spectral measure is supported on the sphere, is a typical model susceptible to fall is this scenario. To obtain the variance asymptotics if the cancellation occurs, one must look at the next non-zero chaotic component and see if it falls in one of the previous aforementioned situations. In any case, the variance asymptotic will be of lower order than the one expected from the heuristic of the case $W=\mathbb{R}$. We refer to Section 1.4 for more details.
1.2 Nodal volume of Gaussian fields
@@
  • For $u,v\in V$ distinct, the vector $(Y(u),Y(v))$ is non-degenerate
@@
It follows from this set of assumption that for $t\in U$, the level set $\{Y=u\}$ is generically a smooth submanifold of codimension $k$ in $V$, to which we associate its random nodal measure $\mathrm{d}Z^{u}$. If $u=0$ we will denote it by $\mathrm{d}Z$. We are interested, in the quantity
@@
Let us study first the projection on the first chaos of $Z_{\lambda}^{u}(\phi)$. By symmetry (see also Lemma 4.5), the first chaotic projection in the case $u=0$ cancels, so the next theorem is relevant for $u\neq 0$. {mdframed}[backgroundcolor=white]
@@
Assume that spectral measure $\psi$ of $Y$ has a continuous density $\Sigma$ w.r.t the Lebesgue measure near the origin. Then
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{u,(1)}(\phi))=\Sigma(0)(u,u).$
If additionally, $r$ and its second differential are square integrable, and $\Sigma(0)(u,u)>0$ then $Z_{\lambda}^{u}(\phi)$ satisfies the CLT.
@@
In the case $u=0$, the two main obstructions in the case $k>1$ to prove general lower bound for the variance by an analysis of the second chaotic component, is that one cannot assume that the random vector $X(0)=(Y(0),\nabla Y(0))$ follows a standard Gaussian distribution by choosing adequate scalar products on $U$ and $V$ (see Remark 4.2), and that it is in general impossible, to compute explicitly the chaotic expansion of the nodal volume without any additional assumption on the field. The computation can be done in particular cases, see Lemma 4.5, under some assumptions of independence. The following theorem proves, under a square integrability property of the covariance function of $Y$, that the variance of the nodal volume converges to a positive constant under classical constraints on the field. {mdframed}[backgroundcolor=white]
@@
Assume that $r$ and its second differential are square integrable on $V$, and one of the three following hypotheses.
  • (H1)
    There are bases of $U$ and $V$ such that $\operatorname{Var}(Y(0),\nabla Y(0))=\mathrm{Id}$
  • (H2)
    There is $u\in U$ such that for all $v\in u^{\perp}$, the process $\langle Y,u\rangle$ and $\langle Y,v\rangle$ are independent
  • (H3)
    $Y$ is a gradient field
Then there is a finite positive constant $C$ such that
@@
This theorem recovers and generalize many previously known results in the literature concerning the positivity of the lower bound for the nodal volume under the square integrability hypothesis.
– Hypothesis $(H1)$ :
This hypothesis assume a strong joint independence assumption on the random variable $(Y(0),\nabla Y(0))$, but assume nothing of the independence of coordinates processes. Following Remark 4.2, this hypothesis is always satisfied if $k=1$.
– Hypothesis $(H2)$ :
This second hypothesis treats the case of the volume of the intersection of the nodal set of a real random field with the nodal set of another independent vector field.
– Hypothesis $(H3)$ :
This last hypothesis proves the variance asymptotics and the CLT for the number of critical points of a random field without any additional assumptions. This hypothesis is independent of the first two because of the non-independence of the second derivatives. It recovers and generalizes some of the recent results concerning critical points such as [Nic17, Aza24].
We were not able to find a satisfying statement that encompasses the three situation, but the heuristic of proof for the positivity of the limiting variance, given in Section 5.2, is that in the three cases we are able to write the density $\Sigma$ of the spectral measure $\mu$ of the process $X=(Y,\nabla Y)$ as a sum of tensor products of two terms: one depending only on the spectral density of the underlying field $\sigma$, and the other one being a polynomial, that comes up from taking the Fourier transform of the derivatives. The conclusion follows by the simple geometric fact that a nonzero square integrable function cannot be supported on the zero set of a non-zero polynomial. In the case where $r\notin L^{2}(V)$, the situation is more intricate and is the content of the next Section.
@@ 1.3 Random waves and second chaos cancellation
We say that $Y$ is a regular random wave if the spectral measure admits a continuously differentiable density w.r.t the measure $\sigma$. The particular case where $\psi=\mathrm{Id}\,\sigma$ (meaning that on an orthonormal basis, $Y$ is a collection of iid isotropic distributions), will be called the isotropic random wave model.
@@
On the torus and the sphere, particularly in dimensions $2$ and $3$, where the spectral decomposition of the Laplace operator is explicit, the model of monochromatic random waves and their nodal set has been intensively studied; see [Ros19, Mar23]. In particular, it has been observed that the variance of various local observables of the isotropic random wave (such as the nodal volume [Wig09, Rud08, Kri13, Mar16, Cam19, Mar20], the defect volume [Mar11], and the Euler characteristic [Cam18b]), in the high-energy limit, is of lower order than one might expect for a "generic" local observable, as suggested by Equation (4.1). This phenomenon is known as the Berry cancellation phenomenon and is purely a mathematical consequence of the second chaos cancellation as $\lambda$ approaches +$\infty$+ ∞. The proofs presented in the aforementioned papers exploit special properties of Bessel functions and the explicit chaos decomposition of the nodal volume functional to establish such general estimates. In the following, we restrict our study to the Euclidean case, which represents the scaling limit of monochromatic random waves on a generic compact manifold. It is expected that this cancellation phenomenon also appears in a broader Riemannian context; see the forthcoming paper [Ste25].
@@
Assume that there are basis of $U$ and $V$ such that $\operatorname{Var}(Y(0),\nabla Y(0))=\mathrm{Id}$. If $Y$ is not a random wave then
@@
In this toy-model example, one explicitly computes the variance of the second chaotic projection to provide a partial equivalence between the inclusion of the support of the spectral measure in the sphere and the second chaos cancellation. In more complicated settings, such as those presented in Theorem 1.8, where explicit computation is not always feasible, one can still establish the positivity of the liminf if the support of the spectral measure is not contained within the zero set of a quadratic form. This approach forms the basis of the proof of Theorem 1.7, under the additional square-integrability assumption. We conjecture that some version of the equivalence given by Theorem 1.8 holds true for critical points, but the computation of the second chaotic component in this context proves to be challenging.
The heuristic, valid for "isotropic" observables such as the nodal volume or some of the aforementioned observables, implies that the isotropic cone of the second chaotic component must form a sphere, whose radius needs to be determined. In the case of the nodal volume, this sphere has a radius precisely $\sqrt{d}/2\pi$. The image of the spectral measure of the field $X=(Y,\nabla Y)$ is a sphere of radius $\sqrt{d}/2\pi$ if and only if $Y$ is a random wave, and the conclusion follows from the second point of Theorem 1.4.
This theorem is a multidimensional generalization of some of the findings in [Lac21], where it was proved in the case $V=\mathbb{R}$ that the limiting variance is positive (possibly infinite) if and only if the process is not a cosine with a Gaussian amplitude. In fact, if the spectral measure, restricted to a domain at a positive distance from the sphere of radius $\sqrt{d}/2\pi$, is not square-integrable, then a straightforward byproduct of the proof, along with Lemma 3.8 and Lemma 3.9, shows that the variance diverges to +$\infty$+ ∞. Together with the discussion after Theorem 1.4, this situation can also lead to a CLT for the quantity $Z_{\lambda}$, inherited from a CLT on $Z_{\lambda}^{(2)}$, provided one can quantify the singularity of the measure $\psi*\psi$ near the origin.
We recover and generalize, for any dimension $d$ and $k$, several of the aforementioned results concerning the second chaos cancellation of the volume of the nodal intersection of $k$ isotropic and independent random waves in $\mathbb{R}^{d}$. For more general $u$-levels, we refer to Remark 5.1. To understand the variance asymptotics of $Z_{\lambda}(\phi)$, one must examine the $4$-th chaos, which is the content of the next Section 1.4.

### v3

@@ Abstract
We establish a general criterion for the positivity of the variance of a chaotic component of local functionals of stationary vector-valued Gaussian fields. This criterion is formulated in terms of the spectral properties of the covariance function, without requiring integrability or isotropy. It offers a simple and robust framework for analyzing variance asymptotics in such models. We apply this approach to the study of the nodal volume and the number of critical points of a Gaussian field, proving the positivity of the limiting variance under mild conditions on the covariance function. Additionally, we examine the asymptotics of nodal volume and critical points of Euclidean random waves, deriving the central limit theorem through an analysis of the second and fourth chaotic components. As a byproduct, we unify and generalize many existing results on the volume of intersections of random waves and their critical points, bypassing the need for traditional, intricate variance computations. Our findings shed new light on the second-chaos cancellation phenomenon from a spectral perspective and can be extended to any local, possibly singular, functional of Gaussian fields.
Acknowledgements
I wish to thank Giovanni Peccati and Michele Stecconi for the numerous discussions about Berry cancellation, and their careful reading of the paper which greatly helped to improve the exposition of the results.
@@ 1.1 A general framework
Let $V$ and $W$ be two finite dimensional vector spaces of dimension $d$ and $n$ respectively. Let $X:V\rightarrow W$ be a vector-valued stationary Gaussian random field with a continuous non-degenerate covariance function $\Omega:V\rightarrow S^{2}(W^{*})$, where this last space denotes the set of symmetric bilinear maps on $W$. We denote by $\eta$ the associated Gaussian measure on $W$, and let $f:W\rightarrow\mathbb{R}$ in $L^{2}(\eta)$. Our primal object of interest is the quantity
@@
A proof of the CLT is usually split into two very distinct problems: the asymptotic normality and the positivity of the limiting variance. The asymptotic normality is well-understood and was first proved in [Cuz76] in the case $d=n=1$ and $\Omega\in L^{1}(\mathbb{R}^{d})$ by approximation using $m$-dependent processes. Since then, it has been refined with the introduction of the Wiener chaos expansion, which has become an indispensable tool for proving asymptotic results related to $Z_{\lambda}(\phi)$. In particular, the celebrated Breuer–Major theorem [Bre83] states that asymptotic normality holds if the Hermite rank of $f$ is $q$ and $\Omega\in L^{q}(V)$. Modern proofs of this theorem now include quantitative CLT and functional convergence [Nou11, Nou20, Nua20] and make use of more advanced techniques, such as Malliavin calculus and Stein’s method. The literature on the topic is extensive.
@@
thus the limiting variance is positive if and only if $\rho(0)$ is positive. In a more general framework, the variance can be understood from local properties of the spectral measure $\rho$ near the origin. To explicitly compute the measure $\rho$ in terms of the spectral measure $\mu$ associated to $X$, the most direct path is to expand $f$ into its orthogonal Hermite decomposition on the space $L^{2}(\eta)$. The variance will then be expressed as sum of powers of the covariance function $\Omega$. Passing to Fourier transform, this yields an expression of $\rho$ in terms of a sum of non-negative terms, given by iterated convolutions of the spectral measure $\mu$ and Hermite expansion of $f$. Additional comments on the advantage of this approach will be made after the statement of Theorem 1.4.
@@ 1.1.1 A general spectral expression for the variance
This corollary, which is a direct corollary of Theorem 1.1 and Lemma 3.8, puts forward the interaction between the iterated convolutions of the spectral measure, their regularity near the origin, and asymptotic results of local functionals of the Gaussian field. In the case $n=1$, the $q$-fold convolution of $\mu$ can be seen as the distribution of a random walk with transition probability distribution $\mu$ after $q$ steps. The properties of $\mu^{*q}$ near the origin are linked to the probability of return around 00 of the random walk. This analogy has been successfully used in the series of papers [Mar08, Mar10], motivated by the probabilistic modeling of the cosmic microwave background radiation, more recently in [Lac22] to obtain variance asymptotics related to the volume of random excursions sets, and in [Gro24] to study the covariance structure of the Berry random wave model.
@@ 1.1.2 A general cancellation theorem for the variance
    If $q$ is even and $\operatorname{Im}(\mu^{*q/2})\subset C(f_{q})$, there is a competition between the cancellation phenomenon and the higher growth of the variance explained in the first two points. From Theorem 1.4, if the inclusion is sufficiently "smooth" then the $q$-th chaos cancels. It might not be the case for more irregular measures, where one can only rely on the general Corollary 1.3. The random waves model, whose spectral measure is supported on the sphere, is a typical model susceptible to fall in this scenario. To obtain the variance asymptotics if the cancellation occurs, one must look at the next non-zero chaotic component and see if it falls in one of the previous aforementioned situations. In any case, the variance asymptotic will be of lower order than the one expected from the heuristic of the case $W=\mathbb{R}$. We refer to Section 1.4 for more details.
1.2 Nodal volume and critical points of Gaussian fields
@@
  • For $v,w\in V$ distinct, the vector $(Y(v),Y(w))$ is non-degenerate
@@
It follows from this set of assumption that for $u\in U$, the level set $\{Y=u\}$ is generically a smooth submanifold of codimension $k$ in $V$, to which we associate its random nodal measure $\mathrm{d}Z^{u}$. If $u=0$ we will denote it by $\mathrm{d}Z$. We are interested, in the quantity
@@
When $U=V$, $u=0$ and $Y=\nabla F$ for a real random field $F$ on $V$, then $dZ$ coincide with the Dirac measure supported on the set of critical points of the random field $F$.
@@
1.2.1 First chaos
Let us study first the projection on the first chaos of $Z_{\lambda}^{u}(\phi)$. By symmetry, the first chaotic projection in the case $u=0$ cancels, so the next theorem is relevant for $u\neq 0$. {mdframed}[backgroundcolor=white]
@@ 1.2.1 First chaos
Assume that $Y(0)\perp\nabla Y(0)$ and that the spectral measure $\psi$ of $Y$ has a continuous density $\Sigma$ w.r.t the Lebesgue measure near the origin. Then there is an explicit positive constant $\alpha_{u}$ such that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{u,(1)}(\phi))=\alpha_{u}\Sigma(0)(u,u).$
If additionally, $r$ is integrable and $\Sigma(0)(u,u)>0$ then $Z_{\lambda}^{u}(\phi)$ satisfies the CLT.
One deduce in particular that the first chaos cancels as soon as the spectral measure is supported on a set disjoint from the origin. This is for instance the case for quantities related to random waves, whose spectral measure is supported on the sphere. This general observation corroborates some findings in [Lac22].
@@
1.2.2 Second chaos
The two main obstructions in the case $k>1$ to prove general lower bound for the variance by an analysis of the second chaotic component, is that one cannot assume that the random vector $X(0)=(Y(0),\nabla Y(0))$ follows a standard Gaussian distribution by choosing adequate scalar products on $U$ and $V$, and that it is in general impossible, to compute explicitly the chaotic expansion of the nodal volume without any additional assumption on the field. The computation can be done in particular cases. We define the following set of hypotheses.
  • (H1)
    There are scalar products on $U$ and $V$ such that $\operatorname{Var}(Y(0),\nabla Y(0))=\mathrm{Id}$.
        This hypothesis assume a strong independence assumption on the random variable $(Y(0),\nabla Y(0))$, but assume nothing of the independence of processes. Up to changing adequately the scalar product on $U$ and $V$, this hypothesis can be assumed to be true if $k=1$.
  • (H2)
    There are scalar products on $U$ and $V$ and $v\in U$ such that for all $w\in v^{\perp}$, $\langle Y,v\rangle\perp\langle Y,w\rangle$
        This hypothesis consider the case of the intersection of the nodal set of a real Gaussian field and the nodal set the nodal set of an another independent Gaussian vector field.
  • (H3)
    $Y$ is the gradient of a real random field $F$.
        Under this hypothesis, the nodal set of $Y$ is exactly the set of critical points of the real random field $F$. This hypothesis is independent of the first two because of the non-independence of the second derivatives and the possible lack of isotropy.
The following theorem proves, under a square integrability property of the covariance function of $Y$, that the variance of the nodal volume converges to a positive constant under one of the following above hypotheses. {mdframed}[backgroundcolor=white]
@@ 1.2.2 Second chaos
Assume that $r$ and its second differential are square integrable on $V$, and one of the three following hypotheses (H1), (H2) or (H3) is satisfied. Then there is a finite positive constant $C$ such that
@@
This theorem recovers and generalize many previously cited results in the literature concerning the positivity of the lower bound for the nodal volume under the square integrability hypothesis.
We were not able to find a satisfying statement that encompasses the three situations, but the heuristic of proof for the positivity of the limiting variance, given in Section 5.2, is that in the two cases we are able to compare the density $\Sigma$ of the spectral measure $\mu$ of the process $X=(Y,\nabla Y)$ to a tensor product of two terms: one depending only on the spectral density of the underlying field $\sigma$, and the other one being a polynomial, that comes up from taking the Fourier transform of the derivatives. The conclusion follows by the simple geometric fact that a nonzero square integrable function cannot be supported on the zero set of a non-zero polynomial. In the case where $r\notin L^{2}(V)$, the situation is more intricate and is the content of the next Section 1.3.
@@ 1.3 Random waves and second chaos cancellation
We say that the random wave $Y$ is
  • regular if its spectral measure admits a continuously differentiable density w.r.t $\sigma$
  • symmetric if its covariance function is symmetric
  • isotropic if $\psi=\mathrm{Id}\,\sigma$, meaning that $Y$ is a collection of iid real-valued random waves.
@@
On the torus and the sphere, particularly in dimensions $2$ and $3$, where the spectral decomposition of the Laplace operator is explicit, the model of monochromatic random waves and their nodal set has been intensively studied; see [Ros19, Mar23]. In particular, it has been observed that the variance of various local observables of the isotropic random wave (such as the nodal volume [Wig09, Rud08, Kri13, Mar16, Cam19, Dal19, Mar20, Not21, Dal21], the defect volume [Mar11], and the Euler characteristic [Cam18b]), in the high-energy limit, is of lower order than one might expect for a "generic" local observable, as suggested by Equation (4.1). This phenomenon is known as the Berry cancellation phenomenon and is purely a mathematical consequence of the second chaos cancellation as $\lambda$ approaches +$\infty$+ ∞. The proofs presented in the aforementioned papers exploit special properties of Bessel functions and the explicit chaos decomposition of the nodal volume functional to establish such general estimates. In the following, we restrict our study to the Euclidean case, which represents the scaling limit of monochromatic random waves on a generic compact manifold. It is expected that this cancellation phenomenon also appears in a broader Riemannian context; see the forthcoming paper [Ste25].
@@
Assume that hypothesis (H1) holds true. If $Y$ is not a random wave then
@@
{mdframed}
[backgroundcolor=white]
Theorem 1.9.
Assume that hypothesis (H3) holds true, and moreover that the field $F$ has an isotropic distribution. If $F$ is not a random wave then
Conversely, if $F$ is the isotropic random wave model, then
[…]
The heuristic of these cancellation phenomena, valid for "isotropic" observables such as the nodal volume or critical points, is that the isotropic cone of the second chaotic component must form a sphere, whose radius needs to be determined. In the case of the nodal observables mentioned above, this sphere has a radius precisely $\sqrt{d}/2\pi$. The image of the spectral measure of the field $X=(Y,\nabla Y)$ is a sphere of radius $\sqrt{d}/2\pi$ if and only if $Y$ is a random wave, and the conclusion follows from the second point of Theorem 1.4.
These theorem are a multidimensional generalization of some of the findings in [Lac21], where it was proved in the case $V=\mathbb{R}$ that the limiting variance is positive (possibly infinite) if and only if the process is not a cosine with a Gaussian amplitude. In fact, if the spectral measure, restricted to a domain at a positive distance from the sphere of radius $\sqrt{d}/2\pi$, is not square-integrable, then a straightforward byproduct of the proof, along with Lemma 3.8 and Lemma 3.9, shows that the variance diverges to +$\infty$+ ∞. Together with the discussion after Theorem 1.4, this situation can also lead to a CLT for the quantity $Z_{\lambda}$, inherited from a CLT on $Z_{\lambda}^{(2)}$, provided one can quantify the singularity of the measure $\psi*\psi$ near the origin.
At last, these theorem recover and generalize, for any dimension $d$ and $k$, several of the aforementioned results concerning the second chaos cancellation of the number of critical points of isotropic random waves, and of the volume of the nodal intersection of $k$ isotropic and independent random waves in $\mathbb{R}^{d}$. For more general $u$-levels, we refer to Remark 5.1.
To understand the variance asymptotics of $Z_{\lambda}(\phi)$, one must examine the $4$-th chaos, which is the content of the next Section 1.4.